# Prompt reinforcing for long-term planning of large language models

### v1

@@ 1 Introduction
Directly optimising LLMs could improve their ability to plan across multiple turns, e.g., supervised fine-tuning with low-rank adaptation (Hu et al., 2022), direct preference optimisation (Feng et al., 2025b), continuous prompting (Lester et al., 2021; Qin & Eisner, 2021; Li & Liang, 2021; Liu et al., 2023), or reinforcement learning with dialogue-level rewards (Feng et al., 2025a); however, these approaches are often impractical for real-time updates due to high computational costs, especially with limited local resources, and are incompatible with API-only LLMs.
@@
Figure 1: The structure of Reinforced Prompt Optimisation (RPO). The initial $\textit{prompt}^{1}$ can be generated by LLMs or written by experts. In interactive optimisation, the system will first interact with the environment, e.g., simulated or real users. The feedbacker, e.g., human experts or LLMs, will provide textual feedback based on trajectories. The rewriter generates a new prompt based on the original prompt and the textual feedback to update the system’s original prompt. One cycle of interactive optimisation is called an epoch, and we use superscripts to denote the epoch number.
@@ 2 Related work
For high parameter counts, training or fine-tuning an entire large language model is infeasible since it requires a huge amount of computational resources. As a result, parameter-efficient fine-tuning, such as training only part of the model or freezing the model and training an adapter, is widely used to refine LLMs (Hu et al., 2022; 2023; Lialin et al., 2023). On the other hand, continuous prompting, e.g., prefix-tuning and soft-prompting, is also popular to adapt LLMs to specific tasks or improve their performance (Lester et al., 2021; Qin & Eisner, 2021; Li & Liang, 2021; Liu et al., 2023). By updating inputs of every attention layer (Li & Liang, 2021), or task-related vectors (Lester et al., 2021), these methods can achieve comparable performance to full fine-tuning across various model sizes and tasks (Liu et al., 2022). Although these methods can improve LLMs effectively, they do not apply to API-access-only LLMs, and such training processes cannot be carried out in real-time.
@@
To improve the performance of text-based prompts, various prompting styles are proposed, e.g., Chain-of-Thought (Wei et al., 2022) or ReAct (Yao et al., 2023). These prompting methods encourage LLMs to reason before taking action or generating responses, which leads to better performance. However, optimising the prompt for better performance by manual trial and error is inefficient. Instead, self-feedback methods are introduced to refine the LLMs’ response, e.g., LLM-augmenter generates feedback by itself and leverages external knowledge to rewrite its response (Peng et al., 2023), and Reflexion summarises previous interactions with the environment as ‘reflections’ to improve the model’s response (Shinn et al., 2023).
@@
Petrov et al. (2024) highlight the limitations of context-based fine-tuning methods, e.g., in-context learning, prompting, and prefix tuning, for new task learning in transformers. Specifically, transformers struggle to acquire new tasks solely through prompting, as prompts cannot change the model’s attention patterns. Instead, they can only bias the output of the attention layers in a fixed direction and elicit skills learned through pre-training. In other words, only models with billions of parameters trained on vast, diverse datasets are capable of in-context learning, adapting to new tasks through examples or instructions without modifying their underlying weights. Therefore, we focus on fundamental models large enough to demonstrate their in-context learning ability, to investigate reinforcement prompt optimisation, which is fully composed of in-context learning with LLMs.
@@ 3 Reinforced Prompt Optimisation
Inspired by the gradient-based optimisation and reinforcement learning algorithms, where a model is initialised from pretraining and then further updated by on-policy learning based on interactions with the environment, we propose the Reinforced Prompt Optimisation (RPO) method (as shown in Figure 1). The initial instruction can be generated by a prompt writer $\textup{LLM}_{P}$ such as the automatic prompt engineer (APE) (Zhou et al., 2023) (the upper left part of Figure 1) or written by human experts (the upper right part of Figure 1).
@@ 3.1 Feedback generation
In contrast, TD-style feedback incorporates turn-level evaluations:
$\textit{feedback}_{\textup{TD},j}=\textup{LLM}_{F}(t_{1},\textit{feedback}_{\textup{TD},1},t_{2},\textit{feedback}_{\textup{TD},2},\dots,t_{j}),$ (2)
where $\textit{feedback}_{\textup{TD},j}$ is the turn-level feedback at turn $j$. All turn-level feedback, $\textit{feedback}_{\textup{TD},1:j}$, will be summarised by $\textup{LLM}_{F}$ into a final $\textit{feedback}_{\textup{TD}}$ afterwards (details of the prompt are shown in Figure 10). Rather than waiting until the dialogue ends, the feedbacker provides incremental assessments at each turn, including the prediction of user sentiment and expected dialogue success, along with actionable suggestions.
@@
where $r_{t}$ corresponds to the short-term reward (e.g., user sentiment after the current turn), $V(s_{t})$ is approximated by the previous turn-level feedback, and $V(s_{t+1})$ represents the estimated long-term value of continuing the dialogue toward successful task completion. This dual perspective enables the system to refine both local decision-making at the turn level and global trajectory planning across the full interaction.
@@ 3.2 Applying Feedback to the Prompt
$\textit{prompt}^{i+1}=\textup{LLM}_{R}(\textit{prompt}^{i},\textit{feedback}^{i},\textit{prompt}^{i-1},\textit{feedback}^{i-1},\ldots,\textit{prompt}^{1},\textit{feedback}^{1}).$ (5)
@@ 4 Experiment Settings
In this study, we focus on iterative meta-prompting by leveraging textual feedback from the environment. We conduct experiments on three challenging human–machine interaction tasks that require multiple turns: Text-to-SQL, Task-oriented Dialogue, and Medical Question-answering (Section 4.1). An overview is shown in Figure 3. Our meta-prompting components are task-agnostic (Section 4.2). They are designed to optimise the prompt of interactive LLM-based systems (Section 4.3). Furthermore, to assess how different prompts affect system performance, all prompts are in a zero-shot in-context learning fashion [^1] Following Brown et al. (2020), this is in-context learning since task descriptions are given as context, but also zero-shot because no demonstrations are included., consisting only of task descriptions without examples.
@@ 4.1 Tasks
To evaluate on a more realistic scenario, we conduct experiments on MultiWOZ 2.1 (Budzianowski et al., 2018; Eric et al., 2020), containing 10k human-to-human conversations on information-seeking, recommendations, and reservations across multiple domains. In this work, we focus on the attraction, hotel, restaurant, and train domains, under the ConvLab-3 framework (Zhu et al., 2023). Each user goal of the simulated user is a plain-text description of requirements, e.g., “You are looking for a place to stay, the hotel should be in the cheap price range and in the city centre. You also need to find a restaurant nearby.”
@@
To evaluate our system in a more human-centred setting and how well prompting can improve the model’s performance in a domain that is not common in the pre-training data, we use two medical question-answering datasets: Huatuo-26M (Wang et al., 2025) and ShenNong-TCM (Wei Zhu & Wang, 2023). The questions in Huatuo-26M and ShenNong-TCM are collected from the internet, e.g., encyclopedias, books, literature, and web corpus, or generated by an LLM based on a traditional Chinese medicine entity graph in Huatuo-26M and ShenNong-TCM, respectively. Simulated users act based on descriptions in plain text, related to general medicine or traditional Chinese medicine, e.g., “我只有咳嗽這一個症狀,請幫我推薦中藥或者方劑。(I only have cough as a symptom. Please recommend Chinese medicine or a prescription.)”.
@@ 4.3 Optimisation and Evaluation
We start by collecting interactions using the initial prompt and user instructions sampled from the training set. The feedbacker receives 10 interactions, since the context length of the LLM-based feedbacker is limited, and to efficiently incorporate human expert feedback. At each epoch, the rewriter generates 2 new prompts based on the previous prompt and the feedback. New interactions are collected with each candidate prompt, and the one with the highest score on the validation set (based on automatic metrics or human experts, depending on the task) is chosen for the next iteration.
@@
In our experiments, we compare three prompt optimisation methods. Automatic Prompt Optimisation (APO) uses the user input, system output, and label to generate feedback (Pryzant et al., 2023). For multi-turn interactions, golden labels are infeasible since multiple solution paths exist; thus, we use a binary success/failure label. Gradient-inspired Prompt Optimizer (GPO) iteratively updates prompts using numerical feedback, e.g., functional accuracy for Text-to-SQL, task success for dialogue (Tang et al., 2025). MC-style (TextGrad) (Yuksekgonul et al., 2025) processes the entire conversation and generates textual feedback, as mentioned in Section 3.1.
@@ 5.1 Robustness and Generalisability
$\text{RPO}_{\text{TD}}$ outperforms prior approaches when the system agent is built with Gemini-2.0-flash, Llama-4-scout, and Llama-3.1-70B. In contrast, $\text{RPO}_{\text{TD+replay}}$ achieves the best overall performance, with an average score of $0.477$ ($54.2\%$ over $\text{Baseline}_{\texttt{Sharded}}$). Llama-3.1-8B benefits the most, since its performance optimised by $\text{RPO}_{\text{TD+replay}}$ ($0.467$) nearly matches the oracle fully-specified setting ($0.505$). The consistent improvements across closed-source (GPT-4o-mini, Gemini-2.0-flash) and open-source (Llama variants) models demonstrate the robustness of our approach and the effectiveness of combining temporal-difference style feedback with replay.
@@
Table 1: Functional accuracy of Text-to-SQL system agents built on five LLMs optimised with various methods. $\text{Oracle}_{\texttt{Full}}$: An oracle baseline in a single-turn setting with fully-specified user queries. The final two columns show the average score (Mean) and the relative improvement ($\Delta\%$) over the $\text{Baseline}_{\texttt{Sharded}}$ across various LLMs. Best scores in the multi-turn setting are bolded.
Method LLM of the system agent Mean $\bf\Delta\%$
GPT Gemini Llama-4 Llama-8B Llama-70B
$\text{Baseline}_{\texttt{Sharded}}$ (Laban et al., 2025) $0.402$ $0.514$ $0.206$ $0.224$ $0.318$ $0.333$ -
APO (Pryzant et al., 2023) $0.374$ $0.523$ $0.318$ $0.290$ $0.336$ $0.368$ $16.9$
GPO (Tang et al., 2025) $0.458$ $0.523$ $0.299$ $0.290$ $0.308$ $0.376$ $17.5$
MC-style (Yuksekgonul et al., 2025) $0.459$ $0.551$ $0.250$ $0.346$ $0.332$ $0.388$ $20.4$
$\text{RPO}_{\text{TD}}$ (ours) $0.439$ $0.561$ $0.336$ $0.318$ $0.383$ $0.408$ $28.9$
$\text{RPO}_{\text{TD+replay}}$ (ours) $\bf 0.528$ $\bf 0.607$ $\bf 0.383$ $\bf 0.467$ $\bf 0.402$ $\bf 0.477$ $\bf 54.2$
$\text{Oracle}_{\texttt{Full}}$ (Laban et al., 2025) $0.893$ $0.841$ $0.729$ $0.505$ $0.748$ $0.743$ $140.2$
@@
Table 2 reports the success rates of FnCTOD  (Li et al., 2024) when optimised by different prompt optimisation methods across five LLM backbones. The baseline system achieves a success rate of $0.420$, while all optimisation methods substantially improve performance. Among prior approaches, MC-style feedback yields the strongest results with a mean success rate of $0.565$ ($34.4\%$ over baseline), slightly outperforming APO and GPO. Our proposed methods consistently surpass these baselines. In particular, $\text{RPO}_{\text{TD}}$ achieves a mean score of $0.575$ ($37.0\%$), demonstrating the advantage of trajectory-driven optimisation. When combined with the rewriter with experience replay, $\text{RPO}_{\text{TD+replay}}$ delivers the best performance across all LLMs, reaching an average success rate of $0.619$, corresponding to a relative improvement of 47.3%. The gains are consistent across all five LLMs, confirming that our approach is robust and generalisable, independent of the underlying model of the meta-prompting agents.
Table 2: The success rate of the task-oriented dialogue system, FnCTOD (Li et al., 2024), improved by various prompt optimisation methods leveraging 5 different LLMs. The initial success rate of FnCTOD is $0.420$. Best scores are bolded.
@@
APO (Pryzant et al., 2023) $0.540$ $0.560$ $0.540$ $0.560$ $0.560$ $0.552$ $31.4$
GPO (Tang et al., 2025) $0.579$ $0.541$ $0.571$ $0.554$ $0.526$ $0.554$ $32.0$
MC-style (Yuksekgonul et al., 2025) $0.567$ $0.549$ $0.575$ $0.560$ $0.572$ $0.565$ $34.4$
$\text{RPO}_{\text{TD}}$ (ours) $0.578$ $0.562$ $0.586$ $0.594$ $0.556$ $0.575$ $37.0$
$\text{RPO}_{\text{TD+replay}}$ (ours) $\bf 0.625$ $\bf 0.622$ $\bf 0.618$ $\bf 0.622$ $\bf 0.606$ $\bf 0.619$ $\bf 47.3$

### v2

@@ 1 Introduction
Directly optimising LLMs could improve their ability to plan across multiple turns, e.g., supervised fine-tuning with low-rank adaptation (Hu et al., 2022), direct preference optimisation (Feng et al., 2025b), continuous prompting (Lester et al., 2021; Qin and Eisner, 2021; Li and Liang, 2021; Liu et al., 2023), or reinforcement learning with dialogue-level rewards (Feng et al., 2025a); however, these approaches are often impractical for real-time updates due to high computational costs, especially with limited local resources, and are incompatible with API-only LLMs.
@@
Figure 1: Structure of Reinforced Prompt Optimisation (RPO). The initial $\textit{prompt}^{1}$ can be written by LLMs or experts. In interactive optimisation, the system will first interact with the environment (simulated or real users). Then, the feedbacker (human or LLM) will provide textual feedback based on trajectories, and the rewriter generates a new prompt based on the original prompt and the textual feedback to update the system’s original prompt. One cycle of interactive optimisation is called an epoch, denoted by superscripts.
@@ 2 Related work
For high parameter counts, training or fine-tuning an entire large language model is infeasible since it requires a huge amount of computational resources. As a result, parameter-efficient fine-tuning, such as training only part of the model or freezing the model and training an adapter, is widely used to refine LLMs (Hu et al., 2022, 2023; Lialin et al., 2023). On the other hand, continuous prompting, e.g., prefix-tuning and soft-prompting, is also popular to adapt LLMs to specific tasks or improve their performance (Lester et al., 2021; Qin and Eisner, 2021; Li and Liang, 2021; Liu et al., 2023). By updating inputs of every attention layer (Li and Liang, 2021), or task-related vectors (Lester et al., 2021), these methods can achieve comparable performance to full fine-tuning across various model sizes and tasks (Liu et al., 2022). Although these methods can improve LLMs effectively, they do not apply to API-access-only LLMs, and such training processes cannot be carried out in real-time.
@@
To improve the performance of text-based prompts, various prompting styles are proposed, e.g., Chain-of-Thought (Wei et al., 2022) or ReAct (Yao et al., 2023). These prompting methods encourage LLMs to reason before taking action or generating responses, which leads to better performance. However, optimising the prompt for better performance by manual trial and error is inefficient. Instead, self-feedback methods are introduced to refine the LLMs’ response, e.g., LLM-augmenter generates feedback by itself and leverages external knowledge to rewrite its response (Peng et al., 2023), and Reflexion summarises previous interactions with the environment as ‘reflections’ to improve the model’s response (Shinn et al., 2023; Madaan et al., 2023).
@@
Petrov et al. (2024) highlight the limitations of context-based fine-tuning methods, e.g., in-context learning, prompting, and prefix tuning, for new task learning in transformers. Specifically, transformers struggle to acquire new tasks solely through prompting, as prompts cannot change the model’s attention patterns. Instead, they can only bias the output of the attention layers in a fixed direction and elicit skills learned through pre-training. Namely, only models with billions of parameters trained on vast, diverse datasets are capable of in-context learning, adapting to new tasks through examples or instructions without modifying their underlying weights. Therefore, we focus on fundamental models large enough to demonstrate their in-context learning ability, to investigate reinforcement prompt optimisation, which is fully composed of in-context learning with LLMs.
@@ 3 Reinforced Prompt Optimisation
Inspired by the gradient-based optimisation and reinforcement learning algorithms, where a model is initialised from pretraining and then further updated by on-policy learning based on interactions with the environment, we propose the Reinforced Prompt Optimisation (RPO) method (as shown in Figure 1 and the pseudo code can be found in Algorithm 1). The initial instruction can be generated by a prompt writer $\textup{LLM}_{P}$ such as the automatic prompt engineer (APE) (Zhou et al., 2023) (the upper left part of Figure 1) or written by human experts (the upper right part of Figure 1).
@@ 3.1 Feedback generation
In contrast, our proposed TD-style feedback incorporates turn-level evaluations:
$\begin{split}\textit{feedback}_{\textup{TD},j}=\textup{LLM}_{F}(&t_{1},\textit{feedback}_{\textup{TD},1},t_{2},\\ &\textit{feedback}_{\textup{TD},2},\dots,t_{j}),\end{split}$ (2)
where $\textit{feedback}_{\textup{TD},j}$ is the turn-level feedback at turn $j$. All turn-level feedback, $\textit{feedback}_{\textup{TD},1:j}$, will be summarised by $\textup{LLM}_{F}$ into a final $\textit{feedback}_{\textup{TD}}$ afterwards (details of the prompt are shown in Figure 19, and examples of turn-level, dialogue-level and final feedback are shown in Figure 9, Figure 10 and Figure 11, respectively). Rather than waiting until the dialogue ends, the feedbacker provides incremental assessments at each turn, including the prediction of user sentiment and expected dialogue success, along with actionable suggestions.
@@
where $r_{t}$ corresponds to the short-term reward (e.g., user sentiment after the current turn), $V(s_{t})$ is approximated by the previous turn-level feedback, and $V(s_{t+1})$ represents the estimated long-term value of continuing the dialogue toward successful task completion [^1] If the task is episodic, the discount factor $\gamma$ can be viewed as set to one (Sutton et al., 2018), so there is no need to specifically take care of it in text.. This dual perspective enables refinement of both local turn-level decision-making and global trajectory planning across the full interaction.
@@ 3.2 Applying Feedback to the Prompt
$\begin{split}\textit{prompt}^{i+1}=&\textup{LLM}_{R}(\textit{prompt}^{i},\textit{feedback}^{i},\textit{prompt}^{i-1},\\ &\textit{feedback}^{i-1},\ldots,\textit{prompt}^{1},\textit{feedback}^{1}).\end{split}$ (5)
@@ 4 Experiment Settings
In this study, we focus on iterative meta-prompting by leveraging textual feedback from the environment. We conduct experiments on three challenging human–machine interaction tasks that require multiple turns: Text-to-SQL, Task-oriented Dialogue, and Medical Question-answering (Section 3). An overview is shown in Figure 3. Our meta-prompting components are task-agnostic (Section 4.2). They are designed to optimise the prompt of interactive LLM-based systems (Section 4.3). Furthermore, to assess how different prompts affect system performance, all prompts are in a zero-shot in-context learning fashion [^2] Following Brown et al. (2020), this is in-context learning since task descriptions are given as context, but also zero-shot because no demonstrations are included., consisting only of task descriptions without examples.
4.1 Tasks
@@
To evaluate on a more realistic scenario, we conduct experiments on MultiWOZ 2.1 (Budzianowski et al., 2018; Eric et al., 2020), containing 10k human-to-human conversations on information-seeking, recommendations, and reservations across multiple domains. In this work, we focus on the attraction, hotel, restaurant, and train domains, under the ConvLab-3 framework (Zhu et al., 2023). Each user goal of the simulated user is a plain-text description, e.g., “You are looking for a place to stay, the hotel should be in the cheap price range and in the city centre. You also need to find a restaurant nearby.”
@@
To evaluate our system in a more human-centred setting and how well prompting can improve the model’s performance in a domain that is not common in the pre-training data, we use two medical question-answering datasets: Huatuo-26M (Wang et al., 2025) and ShenNong-TCM (Wei Zhu and Wang, 2023), which are collected from the internet, e.g., encyclopedias, books, literature, and web corpus, or generated by an LLM based on a traditional Chinese medicine entity graph in Huatuo-26M and ShenNong-TCM, respectively. Simulated users act based on descriptions in plain text, related to general medicine or traditional Chinese medicine, e.g., “我只有咳嗽這一個症狀,請幫我推薦中藥或者方劑。(I only have cough as a symptom. Please recommend Chinese medicine or a prescription.)”.
@@ 4.2 Meta-prompting components
Table 1: Functional accuracy with a 95% confidence interval of Text-to-SQL system agents built on five LLMs optimised with various methods. $\text{Oracle}_{\texttt{Full}}$: An oracle baseline in a single-turn setting with fully-specified user queries. The final two columns show the average score (Mean) and the relative improvement ($\Delta\%$) over the $\text{Baseline}_{\texttt{Sharded}}$. Bold scores are significantly better than the others ($p<0.05$).
Method LLM of the system agent Mean $\bf\Delta\%$
GPT Gemini Llama-4 Llama-8B Llama-70B
$\text{Baseline}_{\texttt{Sharded}}$ $0.402_{\pm 0.002}$ $0.514_{\pm 0.027}$ $0.206_{\pm 0.013}$ $0.224_{\pm 0.016}$ $0.318_{\pm 0.014}$ $0.333$ -
BPO $0.318_{\pm 0.061}$ $0.397_{\pm 0.069}$ $0.220_{\pm 0.026}$ $0.185_{\pm 0.028}$ $0.241_{\pm 0.061}$ $0.272$ 15.7-15.7
APO $0.374_{\pm 0.009}$ $0.523_{\pm 0.008}$ $0.318_{\pm 0.012}$ $0.290_{\pm 0.013}$ $0.336_{\pm 0.004}$ $0.368$ $16.9$
GPO $0.458_{\pm 0.008}$ $0.523_{\pm 0.016}$ $0.299_{\pm 0.017}$ $0.290_{\pm 0.015}$ $0.308_{\pm 0.011}$ $0.376$ $17.5$
MC-style $0.459_{\pm 0.018}$ $0.551_{\pm 0.015}$ $0.250_{\pm 0.013}$ $0.346_{\pm 0.021}$ $0.332_{\pm 0.007}$ $0.388$ $20.4$
$\text{RPO}_{\text{TD}}$ (ours) $0.439_{\pm 0.011}$ $0.561_{\pm 0.013}$ $0.336_{\pm 0.014}$ $0.318_{\pm 0.009}$ $0.383_{\pm 0.011}$ $0.408$ $28.9$
$\text{RPO}_{\text{TD+replay}}$ (ours) $\textbf{0.528}_{\pm 0.011}$ $\textbf{0.607}_{\pm 0.018}$ $\textbf{0.383}_{\pm 0.013}$ $\textbf{0.467}_{\pm 0.022}$ $\textbf{0.402}_{\pm 0.012}$ 0.477 $\bf 54.2$
$\text{Oracle}_{\texttt{Full}}$ $0.893_{\pm 0.007}$ $0.841_{\pm 0.007}$ $0.729_{\pm 0.017}$ $0.505_{\pm 0.017}$ $0.748_{\pm 0.014}$ $0.743$ $140.2$
@@ 4.3 Optimisation and Evaluation
We first collect interactions using the initial prompt and user instructions sampled from the training set. Due to context limits and to facilitate human feedback, the feedbacker processes 10 interactions per epoch. The rewriter then generates 2 new prompts based on the previous prompt and the feedback. New interactions are collected with each candidate prompt, and the one with the highest score on the validation set (based on automatic metrics or human experts, depending on the task) is chosen for the next iteration. The number of training epochs is a manually chosen hyperparameter: we use 5 for Text-to-SQL, 8 for Task-oriented Dialogue, and 3 for Medical Question-Answering.
@@
In our experiments, we compare various prompt optimisation methods. Black-Box Prompt Optimization (BPO) fine-tunes Llama-2-7B-Chat (Touvron et al., 2023) for prompt rewriting based on preference learning (Cheng et al., 2024). Automatic Prompt Optimisation (APO) uses the user input, system output, and label to generate feedback (Pryzant et al., 2023). For multi-turn interactions, golden labels are infeasible since multiple solution paths exist; thus, we use a binary success/failure label. Gradient-inspired Prompt Optimizer (GPO) iteratively updates prompts via numerical feedback, e.g., functional accuracy for Text-to-SQL, task success for dialogue (Tang et al., 2025). MC-style (TextGrad) (Yuksekgonul et al., 2025) processes entire conversations to generates textual feedback (see Section 3.1).
@@ 5.1 Robustness and Generalisability
The prompts optimised by BPO do not improve performance; in some cases, they even degrade it, especially when applied to systems built on different model families, e.g. GPT-4o mini or Gemini-2.0-flash. It is not surprising that BPO is inferior since it is optimised for revising prompts for single-turn tasks and does not generalise well to multi-turn scenarios. Its generalisability is further limited by its dependence on the Llama family as the backend, resulting in suggestions that may not transfer to GPT- or Gemini-based system agents. On the other hand, $\text{RPO}_{\text{TD}}$ outperforms prior approaches when the system agent is built with Gemini-2.0-flash, Llama-4-scout, and Llama-3.1-70B. In contrast, $\text{RPO}_{\text{TD+replay}}$ achieves the best overall performance, with an average score of $0.477$ ($54.2\%$ over $\text{Baseline}_{\texttt{Sharded}}$). Llama-3.1-8B benefits the most, since its performance optimised by $\text{RPO}_{\text{TD+replay}}$ ($0.467$) nearly matches the oracle fully-specified setting ($0.505$). The consistent improvements across closed-source (GPT-4o-mini, Gemini-2.0-flash) and open-source (Llama variants) models demonstrate the robustness of our approach and the effectiveness of combining TD-style feedback with replay.
@@
Table 2 reports the success rates of FnCTOD  (Li et al., 2024) when optimised by different prompt optimisation methods across five LLM backbones. The baseline system achieves a success rate of $0.420$, while all optimisation methods substantially improve performance. Among prior approaches, MC-style feedback yields the strongest results with a mean success rate of $0.565$ ($34.4\%$ over baseline), slightly outperforming APO and GPO. Our proposed methods consistently surpass these baselines. In particular, $\text{RPO}_{\text{TD}}$ achieves a mean score of $0.575$ ($37.0\%$), demonstrating the advantage of trajectory-driven optimisation. When combined with the rewriter with experience replay, $\text{RPO}_{\text{TD+replay}}$ delivers the best performance across all LLMs, reaching an average success rate of $0.619$, corresponding to a relative improvement of 47.3%. The gains are consistent across all five LLMs, confirming that RPO is robust and generalisable, independent of the underlying model of the meta-prompting agents.
Table 2: The success rate with a 95% confidence interval of the task-oriented dialogue system, FnCTOD (Li et al., 2024), improved by various prompt optimisation methods leveraging 5 different LLMs. The initial success rate of FnCTOD is $0.420$. Bold scores are significantly better than the others ($p<0.05$).
@@
APO $0.540_{\pm 0.030}$ $0.560_{\pm 0.018}$ $0.540_{\pm 0.018}$ $0.560_{\pm 0.018}$ $0.560_{\pm 0.018}$ $0.552$ $31.4$
GPO $0.579_{\pm 0.021}$ $0.541_{\pm 0.008}$ $0.571_{\pm 0.046}$ $0.554_{\pm 0.017}$ $0.526_{\pm 0.053}$ $0.554$ $32.0$
MC-style $0.567_{\pm 0.042}$ $0.549_{\pm 0.010}$ $0.575_{\pm 0.038}$ $0.560_{\pm 0.029}$ $0.572_{\pm 0.044}$ $0.565$ $34.4$
$\text{RPO}_{\text{TD}}$ (ours) $0.578_{\pm 0.037}$ $0.562_{\pm 0.036}$ $0.586_{\pm 0.021}$ $0.594_{\pm 0.013}$ $0.556_{\pm 0.042}$ $0.575$ $37.0$
$\text{RPO}_{\text{TD+replay}}$ (ours) $\textbf{0.625}_{\pm 0.038}$ $\textbf{0.622}_{\pm 0.018}$ $\textbf{0.618}_{\pm 0.018}$ $\textbf{0.622}_{\pm 0.007}$ $0.606_{\pm 0.020}$ $\bf 0.619$ $\bf 47.3$
Figure 4: The training curves of different textual feedback generation methods. Each setting uses 4 training seeds and is evaluated on 100 dialogues; the line represents the mean success rate, and the shaded area indicates the standard error.